# Non-equilibrium dynamics of localization phase transition in the non-Hermitian Disorder-Aubry-André model

### v2

@@ IV.2 Numerical results of driven dynamics
In summary, we have conducted a comprehensive study of the static scaling behavior and the driven dynamics of localization transitions in the non-Hermitian DAA model, considering both OBC and PBC. By examining the static behavior of $\xi$, $\rm IPR$ and $\Delta E$, we have demonstrated that the same critical exponents observed under PBC are also valid under OBC. We constructed the KZS for driven dynamics of the non-Hermitian DAA systems and numerically confirmed its validity across three different quench directions. For the dynamical paths we considered, the boundary conditions do not have a significant impact on the dynamical evolution process. Our work generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.

### v3

@@ VI Experimental scheme
In summary, we have conducted a study of the static scaling behavior and the driven dynamics of localization transitions in the non-Hermitian DAA model, considering both OBC and PBC. By examining the static behavior of $\xi$, $\rm IPR$ and $\Delta E$, we have demonstrated that the same critical exponents observed under PBC are also valid under OBC. We constructed the KZS for driven dynamics of the non-Hermitian DAA systems and numerically confirmed its validity across three different local-to-local quench directions. The HKZS in the overlapping critical region of non-Hermitian DAA and Anderson localization has been proposed and numerically confirmed the validity across a local-to-skin quench direction. For the dynamical paths we considered, the boundary conditions do not have a significant impact on the applicability of the KZS. However, during the evolution process of the system, its evolutionary trend will exhibit some differences depending on whether it passes through the skin-effect region. Our work generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.
@@ ACKNOWLEDGMENTS
Figure 13: Scaling properties in the state with the lowest real part of the eigenenergy for fixed $\delta L^{1/\nu_{\delta}}=-0.5$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for 10001000 samples. The double-logarithmic scales are used.
APPENDIX A: NUMERICAL VALIDATION FOR STATIC SCALING EQUATIONS
In this Appendix, we carry out a numerical validation of Eqs. (4) to (6) under OBC with $\delta L^{1/\nu_{\delta}}$ held constant at -0.5. Fig. 13 illustrates the scaling behavior of $\xi$, $\rm IPR$ and $\Delta E$ as functions of $\Delta$. Upon rescaling these quantities in accordance with Eqs. (4) to (6), the resultant curves exhibit a remarkable alignment. This alignment serves as compelling evidence for the validity of these equations.
APPENDIX B: NUMERICAL VALIDATION FOR KZS EQUATIONS
In this Appendix, we numerically verify the KZS scaling equations (Eq. (8) to Eq. (9)) by fixing $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.5$ or $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.5$ under both OBC and PBC respectively. We compute the time evolutions of $\xi$ and $\rm IPR$ for various driving rates $R$, with the initial state with the lowest real part of eigenenergy, and observe that after rescaling according to the KZS equations, the curves for different R converge near the critical region, confirming the validity of these equations as demonstrated in Fig. 14, Fig. 15, Fig. 16, and Fig. 17.
Figure 14: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.5$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
Figure 15: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.5$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
Figure 16: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.5$ under OBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\delta_{0}=1.0$.
Figure 17: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.5$ under PBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\delta_{0}=1.0$.